OpenStudy (anonymous):

Which of the following are solutions to the following equation? 4x^2-100=0 MORE THAN ONE ANSWER A. -5 B. -10 C. 5 D. 10 Which of the following are solutions to the following equation? 4x^2-100=0 MORE THAN ONE ANSWER A. -5 B. -10 C. 5 D. 10 @Mathematics

OpenStudy (anonymous):

x^2=100/4

jimthompson5910 (jim_thompson5910):

4x^2-100=0 4(x^2-25)=0 4(x-5)(x+5)=0 x-5=0 or x+5=0 x=5 or x=-5

jimthompson5910 (jim_thompson5910):

So the answers are choices A and C

OpenStudy (anonymous):

A and C